OpenStudy (anonymous):

if f(x)=13-8x+root(2x^2) and f'(r)=4 find r?

OpenStudy (lgbasallote):

\[f(x) = 13 - 8x + \sqrt{2x^2}\] do you know how to find f(r)?

OpenStudy (anonymous):

no, that's what is confusing me

OpenStudy (lgbasallote):

hint: change all the x with r

OpenStudy (anonymous):

So like this? \[f(x)=13-8\times4+\sqrt{2\times4 ^{2}}\]

OpenStudy (anonymous):

f(r)*

OpenStudy (lgbasallote):

what did you do?

OpenStudy (lgbasallote):

i said replace x with r...not 4

OpenStudy (anonymous):

but since f(r)=4 I thought it was that

OpenStudy (lgbasallote):

no...

OpenStudy (lgbasallote):

just replace x with r

OpenStudy (anonymous):

\[13−8×r+\sqrt{2×r^{2}}\] ?

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

simplify that

OpenStudy (anonymous):

\[13-8r+r \sqrt{2}\]?

OpenStudy (lgbasallote):

yes.

OpenStudy (lgbasallote):

btw...are you sure the question is EXACTLY \[f(x) = 13 - 8x + \sqrt{2x^2}\]

OpenStudy (anonymous):

yes, it's exactly that

OpenStudy (lgbasallote):

you're sure... so \[f(r) = 13 - 8r + r\sqrt 2\] take the derivative of this

OpenStudy (anonymous):

The answer is supposed to be: \[r=3\sqrt{2}\] but the derivative gives me: \[f'(r)=-7\]?